Find three consecutive positive odd integers such that the product of the first and the third is 4 less than 7. Times the second
Answer by ankor@dixie-net.com(22740)

Find three consecutive positive odd integers such that the product of the first and the third is 4 less than 7. Times the second
:
"Find three consecutive positive odd integers"
x, (x+2), (x+4)
:
such that the product of the first and the third is 4 less than 7. Times the second
x(x+4) = 7(x+2) - 4
x^2 + 4x = 7x + 14 - 4
x^2 + 4x = 7x + 10
Combine like terms on the left
x^2 + 4x - 7x - 10 = 0
x^2 - 3x - 10 = 0
Factors to
(x-5)(x+2) = 0
two solutions
x = 5, the positive odd integer, so x=5 is the only solution
:
the three integers, 5, 7, 9
:
:
see if that works in the statement:
"Find three consecutive positive odd integers such that the product of the first and the third is 4 less than 7. Times the second"
5*9 = 7(7) - 4
45 = 49 - 4; confirms our solution